OpenStudy (anonymous):

SAT: If a>0 and 16x^2+bx+9=(4x+a)^2 for all values of x, what is the value of b?

OpenStudy (anonymous):

i am sure there is a snappy way to do this, but lets use a trick instead

OpenStudy (anonymous):

I'm up for anything, haha

OpenStudy (anonymous):

ok i thought i had something snappy,but i guess not lets try this instead, maybe i am being silly it says true for all \(x\) so we know that if \(x=0\) we get \(9=a^2\) and so \(a=3\) or \(a=-3\)

OpenStudy (anonymous):

then \[16x^2+bx+9=(4x+a)^2=16x^2+8ax+a^2=16x^2+8ax+9\] so our two choices are either \(b=8\times 3=24\) or \(b=-24\)

OpenStudy (anonymous):

oooh it says \(a>0\) doh

OpenStudy (anonymous):

oh here are the answer choices: 3 8 16 24 72 Sorry for not posting these up!

OpenStudy (anonymous):

that means \(a=3\) and so \(b=24\)

OpenStudy (anonymous):

choices don't help unless you know how to get them

OpenStudy (anonymous):

You're right. So did you plug a random number in?

OpenStudy (anonymous):

i ignored the \(a>0\) part, this means \(a=3\)

OpenStudy (anonymous):

no, not random

OpenStudy (anonymous):

to solve for \(a\) i chose \(x=0\)

OpenStudy (anonymous):

that way the left hand side is 9 and the right hand side is \(a^2\) then you know \(a=3\)

OpenStudy (anonymous):

So to figure out that a=3, you plugged in zero for x? @satellite73

OpenStudy (anonymous):

yes